An LPG tank contains 11.0 kg of butanr gas (C4H10) in a 20-L gas tank. What mist be the Pressure of the gass inside the tank when the Temperature is 27°C?
1
Expert's answer
2020-12-01T09:13:11-0500

No of moles= mass / molar mass = 11000/58 = 172.4moles

PV=nRT

P = nRT/V

P= 172.4 × 0.0821 × 300 / 20

P= 212.3atm